OpenStudy (anonymous):

A chemist needs 10L of 21% salt solution. The chemist has two salt solutions availible at 15% and 25% salt. Write a solve a linear system to find the volume of each solution that needs to be combined to make the mixture

OpenStudy (anonymous):

let volume of one solution = x volume of other solution=(10-x) (15/100)x + (25/100)(10-x) = (10)(21/100) 0.15x + 2.5 -0.25x=2.1 1x=0.4 x=0.4 L 10-x= 0.6 L

OpenStudy (anonymous):

@brook_27 are u done now

OpenStudy (anonymous):

May God bless you. Thank you ALOT.